# Loss Landscape Degeneracy Drives Stagewise Development in Transformers

### v2

@@ Abstract
Deep learning involves navigating a high-dimensional loss landscape over the neural network parameter space. Over the course of training, complex computational structures form and re-form inside the neural network, leading to shifts in input/output behavior. It is a priority for the science of deep learning to uncover principles governing the development of neural network structure and behavior. Drawing on the framework of singular learning theory, we propose that model development is deeply linked to degeneracy in the local geometry of the loss landscape. We investigate this link by monitoring loss landscape degeneracy throughout training, as quantified by the local learning coefficient, for a transformer language model and an in-context linear regression transformer. We show that training can be divided into distinct periods of change in loss landscape degeneracy, and that these changes in degeneracy coincide with significant changes in the internal computational structure and the input/output behavior of the transformers. This finding underscores the potential of a degeneracy-based perspective for understanding modern deep learning.
[…]
@@ 1 Introduction
  1. 1.
    Transformer training (Section 3): We train two transformers, a language model (LM) with around 3M parameters trained on a subset of the Pile (Gao et al., 2020; Xie et al., 2023), and an in-context linear regression model (LR) with around 50k parameters trained on synthetic regression data following Garg et al. (2022).
  2. 2.
    Degeneracy tracking (Section 4): We quantify loss landscape degeneracy throughout training by estimating the local learning coefficient (LLC; Lau et al., 2025), a measure of degeneracy derived from SLT.
  3. 3.
    Degeneracy-based stage division (Section 5): Motivated by the singular learning process in Bayesian inference Watanabe, 2009, §7.6; Chen et al., 2023, we use critical points in the LLC curve to divide training into developmental stages.
  4. 4.
    Developmental analysis (Sections 6 and 7): We track shifts in each model’s internal computational structure and input/output behavior across training, quantified using various setting-specific metrics.
Crucially, we discover that most of the developmental stages identified by changes in loss landscape degeneracy coincide with significant, interpretable shifts in the internal computational structure and input/output behavior of the transformers, showing that the stage division is meaningful. This discovery represents preliminary evidence of a fundamental link between degeneracy and development in deep learning, underscoring the potential of degeneracy as a lens for understanding modern neural network development. Section 8 discusses this and other implications of our investigation.
@@
$\Delta\hat{\ell}$ $-2.33$ $-1.22$ $-0.18$ $-0.40$ $-0.34$
$\Delta\hat{\lambda}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+26.4}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+22.5}$ ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}-1.57}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+8.62}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+1.77}$
@@
$\Delta\hat{\ell}$ $-0.32$ $-2.21$ $-0.07$ $-0.05$ $-0.029$
$\Delta\hat{\lambda}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+21.4}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+149}$ ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}-12.3}$ ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}-44.1}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+3.56}$
@@
Figure 1: Tracking loss landscape degeneracy reveals developmental stages. We train transformer models on both (a) natural language data and (b) synthetic in-context linear regression data. In addition to test loss (top row), we track loss landscape degeneracy as quantified by the local learning coefficient (LLC) (middle row; Section 4). Critical points in the LLC curve mark boundaries between distinct developmental stages (bottom row; warm hues for increasing LLC, cold for decreasing LLC; Section 5). We show in Sections 6 and 7 that most of these stages coincide with the formation of significant internal structures or changes in input/output behavior. The language model first learns to predict using bigram statistics (LM1), then common $n$-grams (LM2), before forming the induction circuit studied by Olsson et al. (2022) (LM3&LM4). The regression model first learns the optimal context-independent solution (LR1), then acquires robust in-context learning (LR2), then specializes to the pre-training distribution (LR3&LR4). These stage divisions and interpretations are specific to the above training runs, but we show in Section B.4 that similar divisions arise with different training seeds.
@@ 2 Related work
The idea that neural networks development occurs in stages goes back decades (Raijmakers et al., 1996) and has received renewed attention in modern deep learning (e.g., Wei et al., 2022; Olsson et al., 2022; McGrath et al., 2022). In the case of deep linear networks, we understand theoretically that models learn progressively higher-rank approximations of their data distribution (see, e.g., Baldi and Hornik, 1989; Rogers and McClelland, 2004; Saxe et al., 2019) throughout training. Our findings suggest that studying degeneracy could help generalize this understanding to modern architectures that exhibit more complex internal computational structure, such as transformers.
@@
One approach is to visualize low-dimensional slices of the loss landscape (Erhan et al., 2010; Goodfellow et al., 2014; Lipton, 2016; Li et al., 2018; Tikeng Notsawo et al., 2024). Unfortunately, a random slice is with high probability a quadratic form associated to nonzero eigenvalues of the Hessian and is thus biased against geometric features that we know are important, such as degeneracy (Wei et al., 2023). Moreover, Antognini and Sohl-Dickstein (2018) have emphasized the difficulty of probing the loss landscape of neural networks with dimensionality reduction tools.
Other standard methods of quantifying the geometry of the loss landscape, such as via the Hessian, are insensitive to important aspects of degeneracy. For example, the Hessian trace or maximum eigenvalues quantify the curvature of a critical point but ignore degenerate dimensions, and the Hessian rank counts the number of degenerate dimensions but fails to distinguish between dimensions by the order of their degeneracy (e.g., quartic vs. zero). In contrast, the LLC is a principled quantitative measure of loss landscape degeneracy. Section B.5 includes experiments showing that Hessian statistics do not reveal the clear stage boundaries revealed by the LLC in our in-context linear regression setting.
@@ 3 Training transformers in two settings
In this section, we describe both settings and introduce their loss functions and data distributions. Common to both settings is a transformer model denoted $f_{w}$ with parameters $w$, which takes as input a sequence of tokens, also called a context. We describe specific architecture details and training hyperparameters in Sections F.1 and F.2.
@@
Elhage et al. (2021) and Olsson et al. (2022) observed that two-layer attention-only transformers (transformers without MLP layers) form interesting internal computational structures supporting ICL, including induction heads. In order to compare with their behavioral and structural analysis we adopt the same architecture. In Appendix E we also study one-layer attention-only transformers. We note that, while we don’t study language models with MLP layers (following prior work), we do use MLP layers for in-context linear regression.
We consider the standard task of next-token prediction for token sequences taken from a subset of the Pile (Gao et al., 2020; Xie et al., 2023). We denote the input context by $S_{K}=(t_{1},\ldots,t_{K})$ where $K$ is the context length. We denote by $S_{\leq k}$ the prefix context $(t_{1},\ldots,t_{k})$ of context $S_{K}$. Our data is a collection of length-$K$ contexts, $\{S^{i}_{K}\}_{i=1}^{n}$. Thus $S^{i}_{\leq k}$ denotes a prefix of the $i$th context, $S^{i}_{K}$.
@@
The empirical loss is then $\ell_{n}(w)=\frac{1}{K-1}\sum_{k=1}^{K-1}\ell_{n,k}(w)$, with the test loss ^$\hat{\ell}(w)$ defined analogously on a held-out set of examples. The corresponding population loss $\ell(w)$ is defined by taking the expectation with respect to the true distribution of contexts (see also Section A.5).
@@
We consider a standard synthetic in-context linear regression problem. A task is a vector $\mathbf{t}\in\mathbb{R}^{D}$, and an example is a pair $(x,y)\in\mathbb{R}^{D}\times\mathbb{R}$. We sample a context by sampling one task $\mathbf{t}\sim\mathcal{N}(0,I_{D})$ and then sampling $K$ i.i.d. inputs $x_{1},\ldots,x_{K}\sim\mathcal{N}(0,I_{D})$ and outputs $y_{1},\ldots,y_{K}\sim\mathcal{N}(\mathbf{t}^{\top}x,\sigma^{2})$. This results in the context $S_{K}=(x_{1},y_{1},\ldots,x_{K-1},y_{K-1},x_{K})$ with label $y_{K}$. We denote by $S_{\leq k}$ the prefix context $(x_{1},y_{1},\ldots,x_{k})$ of context $S_{K}$, its label is $y_{k}$. Section F.2.2 describes how we encode the $x_{i}$ and $y_{i}$ as tokens. Our data is a set of contexts $\{(\mathbf{t}_{i},S_{K}^{i},y_{K}^{i})\}_{i=1}^{n}$ sampled i.i.d. as described above.
@@
Figure 2: The local learning coefficient (LLC) measures loss landscape degeneracy. The LLC can be defined in terms of the rate at which the parameter space volume (within a given neighborhood and with a given maximum loss) shrinks as the loss threshold is reduced to that of the local minimum. We show four population loss landscapes for a two-dimensional parameter space with decreasing LLC (increasing degeneracy). In these examples, the local multiplicity is 1. See Section A.1 for a detailed description of each example, as well as several additional examples.
@@ 4 Quantifying degeneracy with the local learning coefficient
Given a local minimum ${w^{*}}$ of a population loss $\ell$ (a negative log likelihood), the LLC of ${w^{*}}$, denoted $\lambda({w^{*}})$, is a positive rational number measuring the amount of degeneracy in $\ell$ near $w^{*}$ (Watanabe, 2009; Lau et al., 2025). Intuitively, the loss landscape is more degenerate (lower LLC) at a parameter ${w^{*}}$ if there are more ways in which $w$ can be infinitesimally varied near ${w^{*}}$ such that $\ell(w)$ remains equal to $\ell({w^{*}})$.
[…]
@@
Lau et al. (2025) introduced an estimator for the LLC based on stochastic-gradient Langevin dynamics (SGLD; Welling and Teh, 2011), which we use in our experiments. Let ${w^{*}}$ be a local minimum of the population loss $\ell$. The LLC estimate $\hat{\lambda}({w^{*}})$ is
@@
with inverse temperature $\beta$ (controlling the contribution of the empirical loss landscape) and localization strength $\gamma$ (controlling proximity to ${w^{*}}$). The basic idea behind this estimator is the following: the more degenerate the loss landscape, the easier it is for a sampler exploring the Gibbs posterior to find points of low loss, and, in turn, the lower $\hat{\lambda}({w^{*}})$. Section A.2 discusses technical SGLD details, Section A.3 documents the hyperparameters used in our experiments, and Section A.4 outlines our hyperparameter tuning procedure.
@@
Strictly speaking, the LLC is defined only for loss functions arising as a negative log likelihood, whereas our loss function includes terms from overlapping context prefixes. It is possible to define a negative log likelihood-based loss for transformer training—we show empirically in Section A.5 that this does not have a significant effect on LLC estimates, and so we proceed with overlapping contexts for efficiency.
Moreover, the LLC is defined only for local minima of such loss functions, whereas we note equation 3 is defined for arbitrary ${w^{*}}$ and we apply the estimator throughout training. This approach has precedent in prior work on LLC estimation: Lau et al. (2025) showed that when applied to trained parameters, the estimator accurately recovers the learning coefficient associated with a nearby minimum, and Chen et al. (2023) found that the estimator produces reliable results for parameters throughout training. In our case, we obtain stable estimates throughout training given sufficiently strong localization $\gamma$. See Section A.6 for more details.
Figure 3: In the singular learning process, the Bayesian posterior can shift between neighborhoods with different degeneracy. Watanabe’s free energy formula 4 highlights a tradeoff between loss $\ell_{n}$ (the linear term coefficient) and degeneracy $\lambda$ (the LLC, the logarithmic term coefficient). Consider two local minima $w_{1}^{\ast},w_{2}^{\ast}$ with neighborhoods $W_{1}^{\ast},W_{2}^{\ast}$. As the number of samples $n$ increases, if $w_{2}^{\ast}$ has lower loss and higher LLC than $w_{1}^{\ast}$, $W_{2}^{\ast}$ will suddenly achieve lower free energy than $W_{1}^{\ast}$ at some critical sample size $n_{\text{crit}}$, causing the Bayesian posterior to shift from concentrating in $W_{1}^{\ast}$ to $W_{2}^{\ast}$.

### v3

# Loss Landscape Degeneracy and Stagewise Development in Transformers

@@ Abstract
Deep learning involves navigating a high-dimensional loss landscape over the neural network parameter space. Over the course of training, complex computational structures form and re-form inside the neural network, leading to shifts in input/output behavior. It is a priority for the science of deep learning to uncover principles governing the development of neural network structure and behavior. Drawing on the framework of singular learning theory, we propose that model development is deeply linked to degeneracy in the local geometry of the loss landscape. We investigate this link by monitoring loss landscape degeneracy throughout training, as quantified by the local learning coefficient, for a transformer language model and an in-context linear regression transformer. We show that training can be divided into distinct periods of change in loss landscape degeneracy, and that these changes in degeneracy coincide with significant changes in the internal computational structure and the input/output behavior of the transformers. This finding provides suggestive evidence that degeneracy and development are linked in transformers, underscoring the potential of a degeneracy-based perspective for understanding modern deep learning.
@@ 1 Introduction
$\Delta\hat{\ell}$ 2.33-2.33- 2.33 1.22-1.22- 1.22 0.18-0.18- 0.18 0.40-0.40- 0.40 0.34-0.34- 0.34
[…]
@@
$\Delta\hat{\ell}$ 0.32-0.32- 0.32 2.21-2.21- 2.21 0.07-0.07- 0.07 0.05-0.05- 0.05 0.029-0.029- 0.029
[…]
@@
Figure 1: Tracking loss landscape degeneracy reveals developmental stages. We train transformer models on both (a) natural language data and (b) synthetic in-context linear regression data. In addition to test loss (top row), we track loss landscape degeneracy as quantified by the local learning coefficient (LLC) (middle row; Section˜4). Critical points in the LLC curve mark boundaries between distinct developmental stages (bottom row; warm hues for increasing LLC, cold for decreasing LLC; Section˜5). We show in Sections˜6 and 7 that most of these stages coincide with the formation of significant internal structures or changes in input/output behavior. The language model first learns to predict using bigram statistics (LM1), then common $n$-grams (LM2), before forming the induction circuit studied by Olsson et al. (2022) (LM3&LM4). The regression model first learns the optimal context-independent solution (LR1), then acquires robust in-context learning (LR2), then specializes to the pre-training distribution (LR3&LR4). These stage divisions and interpretations are specific to the above training runs, but we show in Section˜B.4 that similar divisions arise with different training seeds.
  1. 1.
    Transformer training (Section˜3): We train two transformers, a language model (LM) with around 3M parameters trained on a subset of the Pile (Gao et al., 2020; Xie et al., 2023), and an in-context linear regression model (LR) with around 50k parameters trained on synthetic regression data following Garg et al. (2022).
  2. 2.
    Degeneracy tracking (Section˜4): We quantify loss landscape degeneracy throughout training by estimating the local learning coefficient (LLC; Lau et al., 2025), a measure of degeneracy derived from SLT.
  3. 3.
    Degeneracy-based stage division (Section˜5): Motivated by the singular learning process in Bayesian inference Watanabe, 2009, §7.6; Chen et al., 2023, we use critical points in the LLC curve to divide training into approximate developmental stages.
  4. 4.
    Developmental analysis (Sections 6, 7): We track shifts in each model’s internal computational structure and input/output behavior across training, quantified using various setting-specific metrics.
Crucially, we discover that most of the developmental stages identified by changes in loss landscape degeneracy coincide with significant, interpretable shifts in the internal computational structure and input/output behavior of the transformers, showing that the stage division is meaningful. Our investigations are motivated by the hypothesis of a fundamental link between degeneracy and development in deep learning. This hypothesis is theoretically grounded in SLT but so far not empirically validated except in toy models (Chen et al., 2023). We view the above discoveries as preliminary evidence for this hypothesis in larger models, and an indication of the potential of degeneracy as a lens for understanding modern neural network development. Section˜8 discusses this and other implications of our investigation.
@@ 2 Related work
The idea that neural networks development occurs in stages goes back decades (Raijmakers et al., 1996) and has received renewed attention in modern deep learning (e.g., Wei et al., 2022; Olsson et al., 2022; McGrath et al., 2022; Odonnat et al., 2024; Chen et al., 2024; Edelman et al., 2024). In the case of deep linear networks, we understand theoretically that models learn progressively higher-rank approximations of their data distribution (see, e.g., Baldi & Hornik, 1989; Rogers & McClelland, 2004; Saxe et al., 2019) throughout training. Our findings suggest that studying degeneracy could help generalize this understanding to modern architectures that exhibit more complex internal computational structure, such as transformers.
@@
One approach is to visualize low-dimensional slices of the loss landscape (Erhan et al., 2010; Goodfellow et al., 2014; Lipton, 2016; Li et al., 2018; Tikeng Notsawo et al., 2024). Unfortunately, a random slice is with high probability a quadratic form associated to nonzero eigenvalues of the Hessian and is thus biased against geometric features that we know are important, such as degeneracy (Wei et al., 2023). Moreover, Antognini & Sohl-Dickstein (2018) have emphasized the difficulty of probing the loss landscape of neural networks with dimensionality reduction tools.
Other standard methods of quantifying the geometry of the loss landscape, such as via the Hessian, are insensitive to important aspects of degeneracy. For example, the Hessian trace or maximum eigenvalues quantify the curvature of a critical point but ignore degenerate dimensions, and the Hessian rank counts the number of degenerate dimensions but fails to distinguish between dimensions by the order of their degeneracy (e.g., quartic vs. zero). In contrast, the LLC is a principled quantitative measure of loss landscape degeneracy. Section˜B.5 includes experiments showing that Hessian statistics do not reveal the clear stage boundaries revealed by the LLC in our in-context linear regression setting.
@@ 3 Training transformers in two settings
In this section, we describe both settings and introduce their loss functions and data distributions. Common to both settings is a transformer model denoted $f_{w}$ with parameters $w$, which takes as input a sequence of tokens, also called a context. We describe specific architecture details and training hyperparameters in Sections˜F.1 and F.2.
@@
Elhage et al. (2021) and Olsson et al. (2022) observed that two-layer attention-only transformers (transformers without MLP layers) form interesting internal computational structures supporting ICL, including induction heads. In order to compare with their behavioral and structural analysis we adopt the same architecture. In Appendix˜E we also study one-layer attention-only transformers. We note that, while we don’t study language models with MLP layers (following prior work), we do use MLP layers for in-context linear regression.
[…]
@@
The empirical loss is then $\ell_{n}(w)=\frac{1}{K-1}\sum_{k=1}^{K-1}\ell_{n,k}(w)$, with the test loss ^$\hat{\ell}(w)$ defined analogously on a held-out set of examples. The corresponding population loss $\ell(w)$ is defined by taking the expectation with respect to the true distribution of contexts (see also Section˜A.6).
@@
We consider a standard synthetic in-context linear regression problem. A task is a vector $\mathbf{t}\in\mathbb{R}^{D}$, and an example is a pair $(x,y)\in\mathbb{R}^{D}\times\mathbb{R}$. We sample a context by sampling one task $\mathbf{t}\sim\mathcal{N}(0,I_{D})$ and then sampling $K$ i.i.d. inputs $x_{1},\ldots,x_{K}\sim\mathcal{N}(0,I_{D})$ and outputs $y_{1},\ldots,y_{K}\sim\mathcal{N}(\mathbf{t}^{\top}x,\sigma^{2})$. This results in the context $S_{K}=(x_{1},y_{1},\ldots,x_{K-1},y_{K-1},x_{K})$ with label $y_{K}$. We denote by $S_{\leq k}$ the prefix context $(x_{1},y_{1},\ldots,x_{k})$ of context $S_{K}$, its label is $y_{k}$. Section˜F.2.2 describes how we encode the $x_{i}$ and $y_{i}$ as tokens. Our data is a set of contexts $\{(\mathbf{t}_{i},S_{K}^{i},y_{K}^{i})\}_{i=1}^{n}$ sampled i.i.d. as described above.
@@
Figure 2: The local learning coefficient (LLC) measures loss landscape degeneracy. The LLC can be defined in terms of the rate at which the parameter space volume (within a given neighborhood and with a given maximum loss) shrinks as the loss threshold is reduced to that of the local minimum. We show four population loss landscapes for a two-dimensional parameter space with decreasing LLC (increasing degeneracy). In these examples, the local multiplicity is 1. See Section˜A.2 for a detailed description of each example, as well as several additional examples.
@@ 4 Quantifying degeneracy with the local learning coefficient
Given a local minimum ${w^{*}}$ of a population loss $\ell$ (a negative log likelihood), the LLC of ${w^{*}}$, denoted $\lambda({w^{*}})$, is a positive rational number that measures the amount of degeneracy in $\ell$ near $w^{*}$ (Watanabe, 2009; Lau et al., 2025), i.e., how many ways $w$ can be varied near ${w^{*}}$ such that $\ell(w)$ remains equal to $\ell({w^{*}})$. Formally, the LLC is defined as the volume-scaling rate near ${w^{*}}$. This is illustrated in Figure˜2, further described in Section˜A.1, and treated in full detail in Lau et al. (2025). Informally, the LLC is a measure of minimum “flatness.” It improves over conventional (second-order) Hessian-based measures of flatness because the LLC is sensitive to more significant, higher-order contributions to volume-scaling.
@@
Lau et al. (2025) introduced an estimator for the LLC based on stochastic-gradient Langevin dynamics (SGLD; Welling & Teh, 2011), which we use in our experiments. Let ${w^{*}}$ be a local minimum of the population loss $\ell$. The LLC estimate $\hat{\lambda}({w^{*}})$ is
@@
with inverse temperature $\beta$ (controlling the contribution of the empirical loss landscape) and localization strength $\gamma$ (controlling proximity to ${w^{*}}$). The basic idea behind this estimator is the following: the more degenerate the loss landscape, the easier it is for a sampler exploring the Gibbs posterior to find points of low loss, and, in turn, the lower $\hat{\lambda}({w^{*}})$. Section˜A.3 discusses technical SGLD details, Section˜A.4 documents the hyperparameters used in our experiments, and Section˜A.5 outlines our hyperparameter tuning procedure.
@@
Strictly speaking, the LLC is defined only for loss functions arising as a negative log likelihood, whereas our loss function includes terms from overlapping context prefixes. It is possible to define a negative log likelihood-based loss for transformer training—we show empirically in Section˜A.6 that this does not have a significant effect on LLC estimates, and so we proceed with overlapping contexts for efficiency.
Moreover, the LLC is defined only for local minima of such loss functions, whereas we note equation ˜3 is defined for arbitrary ${w^{*}}$ and we apply the estimator throughout training. This approach has precedent in prior work on LLC estimation: Lau et al. (2025) showed that when applied to trained parameters, the estimator accurately recovers the learning coefficient associated with a nearby minimum, and Chen et al. (2023) found that the estimator produces reliable results for parameters throughout training. In our case, we obtain stable estimates throughout training given sufficiently strong localization $\gamma$. See Section˜A.7 for more details.
Figure 3: In the singular learning process, the Bayesian posterior can shift between neighborhoods with different degeneracy. Watanabe’s free energy formula ˜4 highlights a tradeoff between loss $\ell_{n}$ (the linear term coefficient) and degeneracy $\lambda$ (the LLC, the logarithmic term coefficient). Consider two local minima $w_{1}^{\ast},w_{2}^{\ast}$ with neighborhoods $W_{1}^{\ast},W_{2}^{\ast}$. As the number of samples $n$ increases, if $w_{2}^{\ast}$ has lower loss and higher LLC than $w_{1}^{\ast}$, $W_{2}^{\ast}$ will suddenly achieve lower free energy than $W_{1}^{\ast}$ at some critical sample size $n_{\text{crit}}$, causing the Bayesian posterior to shift from concentrating in $W_{1}^{\ast}$ to $W_{2}^{\ast}$.